I would like help on my geometry math problem. It says here that I need to find the measures of two complementary angles if the difference in measure is 18.

Answer by swincher4391(1107)

Let x and y be complementary angles.
Then x-y = 18
y = x+18
By definition, complementary angles add up to 90.
x + y = 90
x + x+18 = 90
x +x = 72
2x = 72
x = 36
y = 36 + 18
y = 54
36,54